$ 16:15: 25 Bias frames and 900-sec Dark

$ 18:45: Sky is mostly clear. Some wispy clouds in the northwest, so it's not photometric.

$ 19:00: Doing Twilight flats in Uspec, V, and g. Paul suggested that pointing 15 degrees above the east horizon would give smaller illumination gradients in the flats, so we're trying that. The dome shutter got stuck trying to open, so we got a late start on flats.

Filter scale counts
Uspec 5 40,000, 30,000, 26,000, 23,000, 20,000
V 6 36,000, 30,000, 23,000, 20,000
g 8 37,000, 33,000, 29,000, 26,000, 22,000

$ 19:40: Focusing at position of GRB20070411

$ 20:16: GRB070411 in SDSS-r after a painful focusing .. first exposure 3 pixels FWHM !!

$ 20:38: GRB070411 second OB, seeing still ~3 pixels. Finished at 21:00 the second OB, stars a bit elongated, seeing ~4 pixels

$ 21:01: Slewing to M81

$ 21:14: M81V - some images are slightly elongated, image quality gradually improves from 1.3 to 0.9 arcsec

$ 21:25: refocus for MS1358

$ 21:29: MS1358-1 - image quality 0.6-0.7 arcsec

$ 22:05: refocus

$ 22:09: MS1358-2 - image quality 0.7 - 0.8 arcsec

$ 22:40: refocus

$ 22:42: MS1358-3 - image quality 0.7 arcsec

$ 23:15: refocus

$ 23:19: MS1358-4 (Uspec) - image quality 0.8 - 0.9 arcsec, poorer in last one

$ 23:58: move to A1835, refocus

$ 00:04: A1835-1 - image quality initially 0.7 arcsec, but degraded to worse than 1 arcsec

$ 00:40: refocus, blurry pupil suggesting seeing has deteriorated significantly

$ 00:43: image quality is about 1 arcsec

$ 01:00: Egami_Lockman_VLA_U_1 - too low to complete block, aborting

$ 01:15: Quasar Point 1 Uspec, focusing on field

$ 01:18: SDT_FAN1_U_1 - Seeing is now 1.5, aborting

$ 01:43: mpobj4_1_g - Seeing is now 1.1 arcsec, going to 0.8 arcsec in the last exposure of the OB Scale was set to 0.3, so the exposure time is too short!!

$ 01:51: mpobj4_1_r - Seeing 0.8 - 0.9 arcsec Scale was set to 0.3, so the exposure time is too short!!

$ 02:01: mpobj4_2_g - Seeing 0.9 arcsec Scale was set to 0.3, so the exposure time is too short!!

090223: Chip4 appears to be missing WCS info 090847: the autoguider was guiding on a galaxy!

$ 02:11: Focusing prior to last OB for mpobj4

$ 02:13: mpobj4_2_r - Seeing 0.7 arcsec Scale was set to 0.3, so the exposure time is too short!!

The seeing has been good on average, but the guider thumbnail is showing some wild swings in seeing...

$ 02:23: focus at position of Quasar Point1 field

$ 02:31: sdt_fan1_U_1 093334: Shutter was apparently closed, because this looks like a Dark

$ 02:42: sdt_fan1_U_1 - Still Dark, aborting Took a test 5-sec R exposure (094701) also Dark.

$ 02:50: Attempting to reboot LBC Filters gave initialization errors: "telnet open error", "Balluff problems on power lines", "Brakes problems on power lines".

$ 02:58: Olga will try restarting using the engineering panel.

$ 03:03: Second attempt to reboot LBC. It seems to have worked.

$ 03:10: Test image - shutter got stuck open this time! (101131)

$ 03:15: Another test image, after Olga did something in the engineering panel. (101520) Hooray! The shutter worked that time. Fingers crossed that it keeps working...

$ 03:18: Uspec focus run - only found one aperture, but correction values were moderate; accepting them

$ 03:33: sdt_fan1_U_1 - Shutter seems to be working (it opened anyway...) Seeing 1.3 - 1.4; the wind is picking up (13 m/s now)

$ 04:05: Since conditions have worsened, we are recomputing the queue, and going to SN2005cf_r1 Seeing seems to be improving, 1.0 arcsec

We seem to have been anti-correlated with the seeing variations tonight; every time we went to the Quasar field, the seeing shot up to >1.3. Then when we went to the poor-seeing fields, it dropped back to <~ 1.0

$ 04:38: Focus run at position of Herc_R.

$ 04:42: Herc_R. We're encroaching on twilight a bit. Seeing is 1 arccsec exposure 1 has background = 10000 exposure 2 has background = 19000 Aborting.

$ 05:05: Twilight flats

Filter scale counts
R 40 20,000, 25,000, 35,000, 43,000, 60,000
B 2.5 10,000, 12,000, 14,000, 17,000, 20,000

-- JasonHarris - 14 Apr 2007
Topic revision: r6 - 15 Apr 2007, JasonHarris
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback